A car is driving around a curve that can be approximated as being circular. What direction does the centripetal force point? O O O O Towards the center of the circle Perpendicular to the plane of the circle In the direction of motion Tangential to the circle Away from the center of the circle The centripetal force (Fc) of an object can be calculated using the following equation: n v where m is the objects mass, v is the objects velocity and R is the radius of the circle. If the radius of thee circle changes by a factor of 0.25, the force changes by a factor of: 4 O 0.8 O 0.3 O 0.25 3 Scroll down. Question continues below.
Answer #1

(a). The centripeatal force act towards the center of the circle.

(b).F C1 /F C2 = r2/r1

i.e., centripeatl force in inversely proportional to radius.

Change in force factor = 1/ 0.25 = 4

(c).Force is directly proportional to square of the velocity.

velocity changes by a factor 4 .So, force changes by a factor od 4 2 = 16